305+ a / 6=66
Thanks! I hope that slash line means divided by...that what I meant. Its been a long time since I've done algebra...so this is stumping me!

Answer by stanbon(48558)

305+ a / 6=66
---------
When you post you should use parentheses so we can tell exactly what you mean.
--------
Meaning #1:
(305+a)/6 66
Multiply both sides by 6 to get:
305+a = 396
a= 396-305= 91
-----------
Meaning # 2:
305 + (a/6) = 66
Multiply both sides by 6 to get:
6*305 + a = 6*66
a= 6*66 - 6*305
a = (66-305)*6
a=-1434
----------
See the difference parentheses makes.
---------
Cheers,
Stan H.